OpenStudy (anonymous):

2w/9 + -8 Solve for w and simplify.. For whatever reason I keep getting 27 and it's wrong.

OpenStudy (anonymous):

we need an equal sign

OpenStudy (anonymous):

Yeah theres one there, I accidentally put + lol. It's 2w/9=-8

jhonyy9 (jhonyy9):

2w/9 =-8 2w=-72 w=-72/2 =-36

OpenStudy (anonymous):

multiply both sides by 9 can you then do it?